OpenStudy (anonymous):

Two equations are shown: Equation A y = −3x − 2 Equation B y equals 3 over x plus 5 Which statement best compares the graphs of the two equations? Both are nonlinear. Both are linear. Equation A is nonlinear and equation B is linear. Equation A is linear and equation B is nonlinear.

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

I think it is both are linear

OpenStudy (anonymous):

Is Equation B \[y=\frac{ 3 }{ x }+5\]or\[y=\frac{ 3 }{ x+5 }\]?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

but I think equation B the line curves a bit

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

Well here's what it looks like on a graph

OpenStudy (anonymous):

What it looks like, very roughly.

OpenStudy (anonymous):

equation b

OpenStudy (anonymous):

Yeah, that's Equation B

OpenStudy (anonymous):

oh so then only equation A is linear

OpenStudy (anonymous):

did you see the kittens on the screen or was that just me

OpenStudy (anonymous):

Mateaus posted it and then deleted it, I suppose

OpenStudy (anonymous):

And yeah, A would be linear but B obviously wouldn't

OpenStudy (anonymous):

can you help me with another question

OpenStudy (anonymous):

Sure, what is it?

OpenStudy (anonymous):

is this a negative or positive association

OpenStudy (anonymous):

As the bottom numbers increase, the left numbers decrease, so a negative association. Source: http://www.ics.uci.edu/~jutts/8/Lecture5Compact.pdf

OpenStudy (anonymous):

using this can you help me figure out wether the basketball team has a higher average height or if there similiar http://learn.flvs.net/webdav/assessment_images/educator_mjmath2_v14/08_14p2_28.jpg

OpenStudy (anonymous):

I got an error when I clicked the link

OpenStudy (anonymous):

I think its there similiar but I'm not sure

OpenStudy (anonymous):

The basketball team is centered further to the right of the graph, so they are going to have a higher average as well.

OpenStudy (anonymous):

hhhmmmm....

OpenStudy (anonymous):

Are we supposed to find the exact difference in the averages?

OpenStudy (anonymous):

I'm not sure the exact question says based on visual inspection

OpenStudy (anonymous):

Ok, so just by looking at the graph we can tell that the basketball players are generally taller.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

There is an 80% chance of rain tomorrow and a 20% chance the football game will be postponed. What is the probability that it will rain tomorrow and the football game will be postponed? 4% 16% 20% 60%

OpenStudy (anonymous):

for this one I thought 60%

OpenStudy (anonymous):

80%=4/5 20%=1/5 \[\frac{ 4 }{ 5 }*\frac{ 1 }{ 5 }=\frac{ 4 }{ 25 }=.16\]

OpenStudy (anonymous):

your suppose to multiply them oh

OpenStudy (anonymous):

Every week William collects a few pounds of plastic bags to recycle. The graph below shows the total number of pounds of plastic bags (y) that William collected in a certain amount of time (x), in weeks: What would most likely be the total amount of plastic, in pounds, William would collect in 10 weeks?

OpenStudy (anonymous):

340 pounds 400 pounds 440 pounds 540 pounds

OpenStudy (anonymous):

I think it be 400 pounds

OpenStudy (anonymous):

40+(n*40) and make n whatever number is the weeks

OpenStudy (anonymous):

or 440 pounds

OpenStudy (anonymous):

the unit rate is 40

OpenStudy (anonymous):

n is the week number in my little formula

OpenStudy (anonymous):

so 40+(10*40) so the answer would be 440

OpenStudy (anonymous):

Yep, that would be it

OpenStudy (anonymous):

Two functions, A and B, are described as follows: Function A y = 8x + 3 Function B The slope is 1 and the y-intercept is 4. How much more is the slope of function A than the slope of function B? 1 7 8 9

OpenStudy (anonymous):

for this one I got 7 but now I'm doubting myself

OpenStudy (anonymous):

Well since 8-1=7, yeah I'd go with 7

OpenStudy (anonymous):

thank you so much I wish I could give you like 5 medals

OpenStudy (anonymous):

Just hit me up with a testimonial if you found my help useful :)